OpenStudy (acespeedfighter):

If x = 2.7, 5x = . (Only input decimals such as 12.7 as the answer.) (1 point) I THINK its 13.5

jimthompson5910 (jim_thompson5910):

You are correct. You replace each 'x' with 2.7 and compute to get 5x = 5*x = 5*2.7 = 13.5